solve 2abs(x-3)+5=15 abs stands for absolute number
solve abs(3x-5)greater and equal to 2

Answer by Nate(3500)

2*abs(x - 3) + 5 = 15
2*abs(x - 3) = 10
abs(x - 3) = 5
x - 3 = +- 5
x = 3 +- 5
x = 8 and x = -2
~~~~~~~~~~~~~~~~~~~~~~~
abs(3x - 5) >= 2
3x - 5 >= +-2
When you cross-a-negative over an inequality sign, you switch it.
3x - 5 >= 2 and 3x - 5 <= -2
3x >= 7 and 3x <= 3
x >= 7/3 and x <= 1
<=====(1]-----[7/3)=====>